OpenStudy (anonymous):

The points A and B have coordinates (7,-1) & (3,3) respectively. Point C is on AB produced such that 2AC=3AB. FInd the coordinates of C. Is the ans (-2.5,5.5)?

OpenStudy (anonymous):

AC = 3/2 * AB \[\large (x_{AC},y_{AC}) = \left(\frac{x_A + k x_B}{1+k},\frac{y_A + k y_B}{1+k} \right) \] k = 3/2 I am not sure but I think it should work.